OpenStudy (anonymous):

Im having problems with this question, I can't figure out the sample variance, here is the information given:

OpenStudy (anonymous):

\[\sum_{i=1}^{1}X^2i = 1,778,400(m^3/ha)^2\] and \[Xbar1= 420m^3 /ha\]

OpenStudy (kirbykirby):

\[ \large \sum_{i=1}^1X_i^2 = 1,778,400\] \[\large \bar{X}=420 \] is that all they gave? is there more data to this?

OpenStudy (anonymous):

here is the original question: The observations (volume in m3/ha ) of ten prism plots from a given forest type are summarized as above Using the information, test if the standard deviation is significantly different from 50 m3/ha.

OpenStudy (kirbykirby):

would that summation be \[\large \sum_{i=1}^{\color{red}{10}} \] ?

OpenStudy (anonymous):

the original was\[\sum_{i=1}^{1}\]

OpenStudy (kirbykirby):

if it was 10 on top, then we could do something, but with just a 1, that is just information from one data point out of 10 given

OpenStudy (anonymous):

okay, perhaps its a typo by my prof. I'm guessing I have to use a test to discover if the two variances are statistically differenct, but I need to know both sample variances first (s^2) is that right?

OpenStudy (kirbykirby):

I think they just want you to test if the variance of the prisms (which you have to calculate) differs statistically from the assumed value of 50.

OpenStudy (kirbykirby):

well standard deviation actually.. not variance

OpenStudy (kirbykirby):

\[\large \begin{align} s^2&=\frac{1}{n-1}\sum_{i=1}^{n}(x_i-\bar{x})^2 \\ &=\frac{1}{n-1}\sum_{i=1}^{n}(x_i^2-2x_i\bar{x}+\bar{x}^2)\\ &=\frac{1}{n-1}\left(\sum_{i=1}^{n}x_i^2-2\bar{x}\sum_{i=1}^{n}x_i+\sum_{i=1}^{n}\bar{x}^2 \right)\\ &=\frac{1}{n-1}\left(\sum_{i=1}^{n}x_i^2-2\bar{x}[n\bar{x}]+n\bar{x}^2\right) \\ &=\frac{1}{n-1}\left(\sum_{i=1}^{n}x_i^2-n\bar{x}^2 \right) \end{align}\] Now here \(n=10\), so \[\large s^2=\frac{1}{10-1}\left(\sum_{i=1}^{n}x_i^2-10\bar{x}^2 \right)=\frac{1}{9}\left(1,778,400-10(420)^2 \right)\] Then the standard deviation, just take the square root of that result.

OpenStudy (anonymous):

Oh I don't think I have ever seen that formula before with the 1/n-1 in front of it.

OpenStudy (kirbykirby):

the population variance, \(\sigma^2\) has 1/n in front. The sample variance, \(s^2\), has 1/(n-1) in front.